# Do LLMs Adhere to Label Definitions? Examining Their Receptivity to External Label Definitions

### v1

@@ 5 Related Work
Label definitions have shown promise in improving zero-shot classification Peskine et al. (2023). While the work establishes the basic utility of definitions, our study delves deeper into how LLMs process and utilize these definitions, particularly in generating explanations. Through our systematic analysis of definition conditions, we reveal crucial insights into whether LLMs truly incorporate external definitions or default to their pre-trained knowledge.
@@ 6 Conclusion and Future Work
Our findings suggest actionable insights for deployment. On edge devices, smaller models (e.g., Phi-3, Mistral) gain the most from carefully tailored definitions in specialized domains, offering lightweight performance boosts. In cloud settings, larger models (e.g., GPT-4) rely more on internal knowledge but often refuse outputs when definitions conflict, making them better suited for high-stakes applications where safety is critical. Even when definitions do not improve classification, they reliably enhance explanation quality, supporting user trust, transparency, and regulatory compliance.
@@
While our study provides strong evidence of the impact of label definitions on LLM performance, certain aspects warrant further exploration. Our analysis focuses on classification tasks, and while the results generalize well within this domain, additional studies in more complex reasoning tasks could provide deeper insights. Additionally, while our experiments rely on prompting strategies, future work could explore fine-tuning methods to further enhance definition adherence. Lastly, although we analyze definition accuracy and alignment, investigating finer-grained model behaviors through interpretability techniques could offer even more precise insights into how LLMs integrate external definitions.
@@ Acknowledgments
We thank Mohammad Eskandari for his valuable assistance in revising the figures in this paper.
@@ Appendix A Additional tables
Few-shot samples Premise: ’A doctor checks on his medical equipment.’ Hypothesis: ’A doctor preparing for work.’ Label: Entailment Premise: ’A child is given a medical check up by a doctor.’ Hypothesis: ’a child with doctor’ Label: Entailment Premise: ’A doctor checks on his medical equipment.’ Hypothesis: ’A doctor getting ready for a patient.’ Label: Neutral Premise: ’A child is given a medical check up by a doctor.’ Hypothesis: ’a sick child with doctor’ Label: Neutral Premise: ’A male doctor looking at a female patients hand.’ Hypothesis: ’The doctor is on a date.’ Label: Contradiction Premise: ’A doctor in blue scrubs performing an surgery.’ Hypothesis: ’A doctor is eating lunch.’ Label: Contradiction
Table 4: Few-shot samples used to generate adjusted definition Appendix F used in the example in Figure 2(c).
Model Version # parameters GPT-4 gpt-4-0613 - LLaMA-3 meta-LLaMA/Meta-LLaMA-3-8B-Instruct 8B Phi-3 microsoft/Phi-3-mini-4k-instruct 3.82B Mistral mistralai/Mistral-7B-Instruct-v0.1 7.24B
@@
Table 6: Template Prompt Used for Mistral in the Definition Scenario, for e-snli. For other scenario prompts, refer to the corresponding code.

### v2

@@ 5 Related Work
Label definitions have shown promise in improving zero-shot classification Peskine et al. (2023). While the work establishes the basic utility of definitions, our study delves deeper into how LLMs process and utilize these definitions, particularly in generating explanations. Through our analysis of definition conditions, we reveal crucial insights into whether LLMs truly incorporate external definitions or default to their pre-trained knowledge.
@@
In summary, our work is unique in four respects: (1) We jointly analyze both knowledge conflict and definition integration, whereas prior work has largely focused on conflict alone; (2) we provide the first cross-domain comparison, revealing that definition receptivity differs substantially between general and specialized domains; (3) we introduce a permutation-based probing strategy that exhaustively tests all possible label–definition mappings, beyond the binary conflict settings previously studied; and (4) we uncover a novel explanation–classification disconnect, showing that gains in explanation quality do not necessarily translate to improved classification accuracy. Together, these contributions differentiate our work from existing literature and open new directions for understanding LLM receptivity to definitions.
@@ 6 Conclusion and Future Work
Our findings suggest actionable insights for deployment. On edge devices, smaller models (e.g., Phi-3, Mistral-7B) gain the most from carefully tailored definitions in specialized domains, offering lightweight performance boosts. In cloud settings, larger models (e.g., GPT-4) rely more on internal knowledge but often refuse outputs when definitions conflict, making them better suited for high-stakes applications where safety is critical. Even when definitions do not improve classification, they reliably enhance explanation quality, supporting user trust, transparency, and regulatory compliance.
@@
An additional avenue for future work is to examine LLM behavior under disputed but valid definitions, where experts may legitimately disagree (e.g., on “fluency” vs. “readability”).
@@
While our study provides strong evidence of the impact of label definitions on LLM performance, certain aspects warrant further exploration. Our analysis focuses on classification tasks, and while the results generalize well within this domain, additional studies in more complex reasoning tasks could provide deeper insights. Additionally, while our experiments rely on prompting strategies, future work could explore fine-tuning methods to further enhance definition adherence. Lastly, although we analyze definition accuracy and alignment, investigating finer-grained model behaviors through interpretability techniques could offer even more precise insights into how LLMs integrate external definitions. We did not employ attention-weight analysis, since prior work has shown that attention weights do not consistently reflect true model decision-making processes (Jain and Wallace, 2019; Wiegreffe and Pinter, 2019; Mohammadi et al., 2024; Naim and Asher, 2024). Nonetheless, we consider probing internal mechanisms, including attention, as valuable future directions.
@@ Acknowledgments
We thank Mohammad Eskandari for his valuable assistance in revising the figures presented in this paper. We also thank the anonymous reviewers for their constructive feedback and suggestions, which helped improve the quality of this work. We gratefully acknowledge support from the UMBC Cybersecurity Leadership – Exploratory Grant and the USISTEF Award. The opinions, conclusions, and recommendations expressed here are solely those of the authors and do not necessarily reflect the views of USISTEF, UMBC, or CrowdStrike.
This material is also based on research that is in part supported by the DARPA for the SciFy program under agreement number HR00112520301. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either express or implied, of DARPA or the U.S. Government.
@@ Appendix A Additional tables
Premise Hypothesis Label A doctor checks on his medical equipment. A doctor preparing for work. Entailment A child is given a medical check up by a doctor. A child with doctor. Entailment A doctor checks on his medical equipment. A doctor getting ready for a patient. Neutral A child is given a medical check up by a doctor. A sick child with doctor. Neutral A male doctor looking at a female patient’s hand. The doctor is on a date. Contradiction A doctor in blue scrubs performing surgery. A doctor is eating lunch. Contradiction
Table 4: Few-shot samples used to generate adjusted definitions (see Appendix F and Figure 2c).
Model Version # parameters GPT-4 gpt-4-0613 - LLaMA-3 meta-LLaMA/Meta-LLaMA-3-8B-Instruct 8B Phi-3 microsoft/Phi-3-mini-4k-instruct 3.82B Mistral-7B mistralai/Mistral-7B-Instruct-v0.1 7.24B
@@
Table 6: Template Prompt Used for Mistral-7B in the Definition Scenario, for e-snli. For other scenario prompts, refer to the corresponding code.